how many quarts of pure antifreeze must be added to 9 quarts of a 30% antifreeze solution obtain a 50% antifreeze solution
Answer by mananth(16946)

percent ---------------- quantity
Alcohol I 30 ---------------- 9 quarts
Alcohol II 100 ---------------- x quarts
Mixture 50 ---------------- 9 + x quarts

30*9+100x=50(9+x)
270 + 100 x = 450 + 50 x
100 x + -50 x = 450 - -270
50 x = 180
/ 50
x = 3.6
3.6 quarts of 100 % Alcohol II